# Imputation of Missing Values in Multi-View Data

### v3

@@ 6 Discussion
We proposed a meta-level imputation method for missing values in multi-view data. We evaluated the proposed method with several imputation algorithms using simulations and an application to real multi-view MRI data. We have compared the meta-level implementation of these imputation algorithms with their feature-level implementation whenever possible, and also compared the results with those of a dedicated multi-view imputation method (MOFA), as well as complete case analysis and complete data analysis. In line with Orchard and Woodbury [86, p. 697], who stated that “the best way to treat missing data is not to have them”, we found that complete data analysis often performs best. Often, but not always: In a real multi-view data set, if we removed 75% of observed values for 31 out of the 40 views and imputed them using the unconditional mean, predictive accuracy actually improved slightly. At a first glance it seems counter-intuitive that removing large volumes of data would improve performance. However, recall that MI reduces the variance of imputed features and attenuates the correlation with the outcome [18]. As such, mean imputation is a kind of regularization, applied only to the views in which missing values occur. If a view is very noisy, this regularization can be beneficial, but if the view contains strong signal, this regularization can be detrimental. Note, however, that since we use models that already contain explicit regularization terms, the exact effect on the performance of the model will depend on the interplay between the view-specific regularization caused by mean imputation and the overall regularization through the penalty parameters. Nevertheless, it is clear that the performance of mean imputation heavily depends on whether the missing data corresponds to signal or to noise. Obviously, whether a view corresponds to signal or to noise is generally unknown beforehand. Thus, unconditional mean imputation cannot be recommended as a general imputation method, and is especially unsuitable if the goal is not just to predict, but also to perform view selection.
@@
Our results show that meta-level imputation is much faster than feature-level imputation, at no or very minor costs to predictive and view-selection performance. In fact, meta-level imputation allows the use of sophisticated imputation algorithms such as predictive mean matching (PMM) or missForest (MF) which are computationally heavy or infeasible when applied at the feature level. Meta-level imputation using PMM or MF performed better than MOFA-based imputation.
@@
For the meta-level implementation of PMM, we used two different ways of combining the imputed data sets. The first one is generating the $\bm{Z}$ matrix once based on observed data, imputing it 5 times, then training the meta-learner on the averaged data (mPMM). The second is generating the $\bm{Z}$ matrix 5 times, imputing each of these matrices once, and then training the meta-learner on the averaged data (cvPMM). The cvPMM approach is much more computationally expensive (Figures 8 and 7), but also averages over different allocations to the cross-validation folds, instead of just over different imputations. However, in our simulations, this approach did not lead to an improvement in classification performance, and only small differences in view selection performance. In our real data application we only applied the mPMM variant, which performed considerably worse than mMF. It is possible that cvPMM would have performed better, but given the results of our simulations we found this to be unlikely and not worth the additional computation cost. Of course, other methods of combining imputations are possible. For example, one could train the meta learner separately on each imputated data set, leading to 5 separate StaPLR models. However, this leads to additional questions in how to combine the separate models, and given the observed differences between mPMM and cvPMM, we would not expect this approach to cause large gains in performance.
@@
Using missForest in our simulations, we found that feature-level MF sometimes performed better than mMF or mPMM/cvPMM in terms of MSEP, but these observed differences in terms of predicted probabilities did not affect classification performance. In our real data analysis, mMF was also the best-performing imputation method in terms of classification accuracy (excluding the aforementioned use of mean imputation for the functional views). For view selection, the observed differences between MF and mMF/mPMM/cvPMM are not entirely consistent. However, it is worth noting that in some cases base level MF lead to considerably higher values of FPR and FDR than mMF. Thus, overall it can be said that the meta-level implementation of missForest does not perform worse than the base-level implementation in most important metrics. Additionally, in the real data the coefficients obtained from training the meta-learner on the values imputed by mMF were the closest to those obtained from the complete data. Thus, out of all the imputation methods for multi-view data that we considered in this study, we would most recommend meta-level missForest imputation because it consistently performed well in both the simulations and real data application, and is much faster than the feature-level implementation.
@@
Throughout this article we have assumed that while the number of features $m$ was larger than $n$, the number of views $V$ was always smaller. It is natural then to wonder how the meta-level imputation methods would perform if $n<V$ (i.e. if $\bm{Z}$ were high-dimensional). Given that missForest can handle high-dimensional data, as is evident by the performance of MF in this article, mMF could most likely be used without issue. The same cannot be said for mPMM or cvPMM, since standard PMM based on the linear model struggles even with small high-dimensional data sets (A). However, alternative approaches could be constructed. For example, an imputation method based on the lasso [87, 88] is included in mice [23], and appeared to work quite well for a small high-dimensional data set such as the one described in A (results not shown), although the associated computation time still made it infeasible to include as a base-level imputation method in this study. Nevertheless, since generally $V\ll m$, it has potential as a meta-level imputation method in cases where $\bm{Z}$ is high-dimensional.
@@
In our simulations missingness occurred only in the training set, but in practical settings it is likely that missingness occurs also in the test set. In this case, Josse et al. [19] recommend that the same imputation model that was used to impute the training set is used to impute the test set. Unfortunately, this is not so easy to do in practice, since many implementations of imputation methods, including mice [23] and missForest [28], effectively work as ‘black boxes’ that do not separate the imputation model from its use to complete the data [19]. A possible solution that has been suggested for this problem is to take a semi-supervised approach [19], jointly imputing the train and test set [89], and then learning the prediction model on the training set only, but this does require that the train and test set are available at the same time [19].
@@ Appendix A Infeasibility of Predictive Mean Matching Using MICE
Even when good results could be obtained, two practical problems make calculations using mice infeasible for large high-dimensional problems. The first, and most important one, is time: Even for the simplest setting included in our experiments, a single iteration of a single imputation of a single variable took well over 24 hours to compute. The second, which was not directly relevant to our study but will be for larger problems, is memory handling: A test run of PMM with mice on a data set consisting of 200 observations of 75,000 variables (120 MB of data), which is small by medical imaging standards, already required 42 GB of RAM.

### v4

@@ 6 Discussion
We proposed a meta-level imputation method for missing values in multi-view data. We evaluated the proposed method with several imputation algorithms using simulations and an application to real multi-view MRI data. We compared the meta-level implementation of these imputation algorithms with their feature-level implementation whenever possible, and also compared the results with those of a dedicated multi-view imputation method (MOFA), as well as complete case analysis and complete data analysis. In line with Orchard and Woodbury [86, p. 697], who stated that “the best way to treat missing data is not to have them”, we found that complete data analysis often performs best. Often, but not always: In a real multi-view data set, if we removed 75% of observed values for 31 out of the 40 views and imputed them using the unconditional mean, predictive accuracy actually improved slightly. At a first glance it seems counter-intuitive that removing large volumes of data would improve performance. However, recall that MI reduces the variance of imputed features and attenuates the correlation with the outcome [18]. As such, mean imputation is a kind of regularization, applied only to the views in which missing values occur. If a view is very noisy, this regularization can be beneficial, but if the view contains strong signal, this regularization can be detrimental. Note, however, that since we use models that already contain explicit regularization terms, the exact effect on the performance of the model will depend on the interplay between the view-specific regularization caused by mean imputation and the overall regularization through the penalty parameters. Nevertheless, it is clear that the performance of mean imputation heavily depends on whether the missing data corresponds to signal or to noise. Obviously, whether a view corresponds to signal or to noise is generally unknown beforehand. Thus, unconditional mean imputation cannot be recommended as a general imputation method, and is especially unsuitable if the goal is not just to predict, but also to perform view selection.
@@
Our results show that meta-level imputation is much faster than feature-level imputation, at no or very minor costs to predictive and view-selection performance. In fact, meta-level imputation allows the use of sophisticated imputation algorithms such as predictive mean matching (PMM) or missForest (MF) which are computationally heavy or infeasible when applied at the feature level. In our experiments meta-level imputation using PMM or MF performed better than MOFA-based imputation.
@@
For the meta-level implementation of PMM, we used two different ways of combining the imputed data sets. The first one is generating the $\bm{Z}$ matrix once based on observed data, imputing it 5 times, then training the meta-learner on the averaged data (mPMM). The second is generating the $\bm{Z}$ matrix 5 times, imputing each of these matrices once, and then training the meta-learner on the averaged data (cvPMM). The cvPMM approach is much more computationally expensive (Figures 8 and 7), but also averages over different allocations to the cross-validation folds, instead of just over different imputations. However, in our simulations, this approach did not result in an improvement in classification performance, and only small differences in view selection performance. In our real data application we only applied the mPMM variant, which performed considerably worse than mMF. It is possible that cvPMM would have performed better, but given the results of our simulations we found this to be unlikely and not worth the additional computation cost. Of course, other methods of combining imputations are possible. For example, one could train the meta learner separately on each imputated data set, leading to 5 separate StaPLR models. However, this leads to additional questions on how to combine the separate models, and given the observed differences between mPMM and cvPMM, we would not expect this approach to cause large gains in performance.
@@
Using missForest in our simulations, we found that feature-level MF sometimes performed better than mMF or mPMM/cvPMM in terms of MSEP, but these observed differences in terms of predicted probabilities did not affect classification performance. In our real data analysis, mMF was also the best-performing imputation method in terms of classification accuracy (excluding the aforementioned use of mean imputation for the functional views). For view selection, the observed differences between MF and mMF/mPMM/cvPMM are not entirely consistent. However, it is worth noting that in some cases base level MF resulted in considerably higher values of FPR and FDR than mMF. Overall, the meta-level implementation of missForest does not perform worse than the base-level implementation in most important metrics. Additionally, in the real data the coefficients obtained from training the meta-learner on the values imputed by mMF were the closest to those obtained from the complete data. Thus, out of all the imputation methods for multi-view data that we considered in this study, we would most recommend meta-level missForest imputation because it consistently performed well in both the simulations and real data application, and is much faster than the feature-level implementation.
@@
Throughout this article we have assumed that while the number of features $m$ was larger than $n$, the number of views $V$ was always smaller. It is natural then to wonder how the meta-level imputation methods would perform if $n<V$ (i.e., if $\bm{Z}$ were high-dimensional). Given that missForest can handle high-dimensional data, as is evident by the performance of MF in this article, mMF could most likely be used without issue. The same cannot be said for mPMM or cvPMM, since standard PMM based on the linear model struggles even with small high-dimensional data sets (see A). However, alternative approaches could be constructed. For example, an imputation method based on the lasso [87, 88] is included in mice [23], and appeared to work quite well for a small high-dimensional data set such as the one described in A (results not shown), although the associated computation time still made it infeasible to include as a base-level imputation method in this study. Nevertheless, since generally $V\ll m$, it has potential as a meta-level imputation method in cases where $\bm{Z}$ is high-dimensional.
@@
In our simulations missingness occurred only in the training set, but in practical settings it is likely that missingness occurs also in the test set. In this case, Josse et al. [19] recommend that the same imputation model that was used to impute the training set is used to impute the test set. Unfortunately, this is not easy to do in practice, since many implementations of imputation methods, including mice [23] and missForest [28], effectively work as ‘black boxes’ that do not separate the imputation model from its use to complete the data [19]. A possible solution that has been suggested for this problem is to take a semi-supervised approach [19], jointly imputing the training and test set [89], and then learning the prediction model on the training set only, but this does require that the training and test set are available at the same time [19].
@@
This article focuses on the imputation of missing values. As previously stated in Section 2, other methods for handling missing data exist which do not explicitly impute missing values, including likelihood-based methods such as full information maximum likelihood (FIML) [34, 35]. Such methods may provide an elegant way for dealing with missing values in models based on maximum likelihood estimation, or extensions thereof such as multi-view (deep) Gaussian process modeling [90, 91]. However, such methods can have numerous computational difficulties since the resulting optimization problem typically has no closed-form solution and may have many local optima [92, 93]. Multiple imputation methods are more broadly applicable [18, 19, 35], and have the added advantage of making the imputed values explicit, allowing them to be inspected and analyzed [18].
@@ Appendix A Infeasibility of Predictive Mean Matching Using MICE
Even when good results could be obtained, two practical problems make calculations using mice infeasible for large high-dimensional data sets. The first, and most important one, is time: Even for the simplest setting included in our experiments, a single iteration of a single imputation of a single variable took well over 24 hours to compute. The second, which was not directly relevant to our simulation study but will be for larger problems, is memory handling: A test run of PMM with mice on a data set consisting of 200 observations of 75,000 variables (120 MB of data), which is small by medical imaging standards, already required 42 GB of RAM.